Problem3 Leon Company makes electronic equipment. Each unit is assembled by hand It is checked for quality before being released for sale This is some of the data related to operations for a year: Sales Price $ 350.00 Variable Costs (per unit sold) Assembly Labor Manufacturing overhead (MOH) Quality control labor Sales Commission Materials 50.00 20.00 10.00 2.50 2.50 Fixed Costs (per year) Manufacturing overhead Sales & Administration 112,500 125,400 Calculate the amount of each of the following if 1,200 units are produced and if 1,000 units are produced during a year (assuming the activity is within the relevant range) Production Level 1,200 1,000 Variable manufacturing cost per unit Total variable manufacturing cost per year Fixed manufacturing cost per unit Total fixed manufacturing cost per year Total manufacturing cost per unit Total nonmanufacturing cost per unit
Answer #1

Calculate following

Production level
1200 1000
Variable manufacturing cost per unit (50+20+10+2.50) = 82.50 82.50
Total variable manufacturing cost per year 82.50*1200 = 99000 82.50*1000 = 82500
Fixed manufacturing cost per unit 112500/1200 = 93.75 112500/1000 = 112.50
Total fixed manufacturing cost per year 112500 112500
Total manufacturing cost per unit 93.75+82.50 = 176.25 82.50+112.50 = 195
Total non manufacturing cost per unit (125400/1200+2.50) = $107 (125400/1000+2.50) = $127.90